# Xingwang Li, , Dusit Niyato, and Chau Yuen

### v1

@@ IV-B FIM Channel Estimation as Neural Operator Learning
A naive application of generic black-box models is ill-suited for this problem, as they lack the necessary inductive biases to handle the continuous deformation space, disregard the underlying wave physics, and suffer from mesh-dependence, making them impractical and data-inefficient [33]. In this work, we reformulate the FIM channel estimation problem as learning an operator $\mathcal{G}$ that maps the channel response function at a given set of pilot deformations $\{\bm{\zeta}_{m}\}_{m=1}^{M}$, along with a query for a target deformation, to the channel vector at that target deformation. The goal is to learn the operator $\mathcal{G}$ such that for any arbitrary and continuous target deformation $\mathbf{\bm{\zeta}}_{\text{target}}$, the corresponding channel vector estimate $\hat{\mathbf{h}}(\mathbf{\bm{\zeta}}_{\text{target}})$ can be expressed as
@@ IV-B1 Lifting Layer
The input to the FNO is a function defined on the spatial domain of the FIM, which is discretized by $N$ antenna elements. For each antenna element $n$, we construct a high-dimensional feature vector $\mathbf{v}_{\text{in}}(n)\in\mathbb{R}^{3M+1}$ that encapsulates all available information: the measured channel responses and the FIM shape information. This feature vector is defined as
@@ IV-B2 Fourier Layers
where $\sigma$ is a non-linear activation function. The update of the Fourier layer is composed of two parallel paths: local transformation $\mathcal{W}_{t}\mathbf{v}_{t}(n)$ and global kernel integration $\left(\mathcal{K}(\mathbf{v}_{t};\mathbf{\varphi}_{t})\right)(n)$. The term $\mathcal{W}_{t}\mathbf{v}_{t}(n)$ is a pointwise linear transformation applied at each location $n$ with learnable weight $\mathcal{W}_{t}$, typically implemented as a 1D convolution with a 1$\times$1 kernel. It captures local and channel-wise relationships. The term $\left(\mathcal{K}(\mathbf{v}_{t};\mathbf{\varphi}_{t})\right)(n)$ is equivalent to a pointwise multiplication in the Fourier domain:
@@
Remark 3: The Fourier-domain operation allows the FNO to efficiently learn global dependencies across the entire FIM array with a constant number of parameters, irrespective of the number of antennas $N$. This grants the FNO a powerful property of mesh-independence. A model trained on an FIM with $N$ antennas can be directly evaluated on a large-scale FIM with $N^{\prime}>N$ antennas without retraining. This offers immense practical value for deploying solutions across FIM hardware of varying scales and configurations.
@@ IV-B3 Projection Layer
[$\begin{split}\left[\Re\left\{\hat{\mathbf{h}}(\mathbf{\bm{\zeta}}_{\text{target}})_{n}\right\},\Im\left\{\hat{\mathbf{h}}(\mathbf{\bm{\zeta}}_{\text{target}})_{n}\right\}\right]^{T}=\mathcal{Q}\left(\mathbf{v}_{T}(n)\right).\end{split}$ (42)
@@ IV-C1 Encoder Path
The output is then downsampled by an average pooling operator $\mathcal{A}$, i.e., $\mathbf{v}_{l}=\mathcal{A}(\mathbf{v}^{\prime}_{l})$. This process is repeated for $L_{\text{enc}}$ levels, resulting in a sequence of latent representations $\{\mathbf{v}^{\prime}_{1},\mathbf{v}^{\prime}_{2},\dots,\mathbf{v}^{\prime}_{L_{\text{enc}}}\}$ with progressively decreasing spatial dimensions ($N,N/2,\dots,N/2^{L_{\text{enc}}-1}$).
@@ IV-C3 Decoder Path
[…]
@@
As illustrated in Section IV-A, the proposed H-FNO $\mathcal{G}_{\Theta}$ can be interpreted as a learned approximation to the MAP estimator for the FIM channel estimation problem, where the underlying prior imposed by the H-FNO architecture aligns with the inherent spectral properties of the physical channel model. The kernel integration in the Fourier domain Eqs. (39)-(41) of the H-FNO is mathematically equivalent to placing a Gaussian process (GP) prior on the channel function space. For the data fidelity term of MAP, i.e., the first term in Eq. (34), the H-FNO is trained end-to-end by minimizing the MSE loss in Eq. (46), which is a direct data-driven approximation of the data fidelity term. For the second regularization term in Eq. (34) that penalizes high-frequency channel content, the H-FNO implements this regularization structurally. The truncation of Fourier modes beyond $\kappa_{\text{max}}$ is a hard constraint, effectively setting the prior probability of very high-frequency functions to zero. Within the preserved modes, the learnable transformation $\mathbf{R}_{\varphi}$ can be interpreted as learning the inherent structure of the channel function space.
@@
Remark 4: In the popular deep learning-enabled channel estimators, the deep neural network does not require an explicit prior, which implicitly imposes a learned structural prior on the function space of possible channels as a generic function approximator. However, the proposed H-FNO is not a closed box but a principled and data-driven method for finding a powerful approximation $\mathcal{G}_{\Theta}$, to the true MAP estimator. Its architecture is explicitly designed to efficiently search within a function space constrained by a physically-motivated GP prior. This theoretical synergy between the inductive biases of the proposed H-FNO and the physical characteristics of FIM provides a rigorous justification for its application.
@@ V Numerical Results
V-A Parameter Setups an Dataset Construction
In the simulations, we consider the BS is located at a fixed position of (0, 0, 10) m, while the $K$ users are randomly and uniformly distributed within a rectangular volume defined by the diagonal corner coordinates of (-10, 10, 0) m and (10, 30, 0) m. Unless otherwise specified, we set the FIM size to $N=N_{x}\times N_{z}=8\times 8$, the number of pilot shapes to $M=16$, the number of users to $K=4$, the maximum deformation range to $\bar{\zeta}=0.5\lambda$, the number of scattering clusters to $L=5$ with $G=6$ paths per cluster, the path loss exponent to $v=2.2$, and the carrier frequency to $f_{c}=28$ GHz. The performance of all algorithms is evaluated using the normalized mean squared error (NMSE), defined as $\text{NMSE}=\mathbb{E}\{{{||\hat{\mathbf{h}}_{k}(\bm{\zeta_{\text{target}}})-\mathbf{h}_{k}(\bm{\zeta_{\text{target}}})||}_{F}^{2}/{||\mathbf{h}_{k}(\bm{\zeta}_{\text{target}})||_{F}^{2}}}\}$.
In the training stage of the proposed H-FNO, a dataset of $N_{s}=20,000$ samples is synthetically generated, with $N_{v}=1000$ of these samples being reserved for the validation set. For each sample, a unique channel environment is realized based on the clustered multipath model, and then $M$ noisy pilot observations are generated, where the corresponding noise-free channel vector $\mathbf{h}_{k}(\bm{\zeta}_{\text{target}})$ serves as the ground-truth label. To enhance the model robustness, the SNR for each training batch is randomly selected from {0, …, 20} dB. The model is trained for $E_{s}=50$ epochs with a batch size of $B_{s}=64$. For each SNR point under performance testing, the NMSE results are averaged over $U=500$ Monte Carlo simulations. In each simulation trial, both a new channel realization and a new random target deformation shape are generated to ensure a robust and fair evaluation. For the model-based algorithms, the number of nearest neighbors for KNN is set to $\bar{k}=5$, and the OMP dictionary is constructed from a discrete grid of $D=256$ angles. For the H-FNO, we set the latent dimension to $d_{v}=64$ and the number of encoder levels to $L_{\text{enc}}=2$, with the number of Fourier modes per level set as $\kappa_{l,\text{max}}=16/2^{l}$, which represent a standard architecture designed to balance model expressiveness with computational efficiency.
@@ V-B Performance Comparison between Different Algorithms
Fig. 5 presents the NMSE performance of different channel estimation algorithms versus the signal-to-noise ratio (SNR), where the channel estimation accuracy of all methods improves with increasing SNR. Among the conventional model-based methods, OMP shows strong performance, benefiting from the inherent sparsity of the mmWave channel, while interpolation-based methods like KRR and KNN provide a stable but less accurate estimation. The proposed H-FNO model consistently and significantly outperforms all other conventional model-based algorithms across the entire SNR range, demonstrating its powerful capability in learning the complex mapping from FIM shapes to channel responses and its robustness against noise.
@@ V-C Generalization Performance of the Proposed H-FNO
Figure 7: NMSE performance of different number of antennas $N$ for the proposed H-FNO.
@@
Fig. 8 validates the zero-shot generalization capability of the proposed H-FNO model across different FIM hardware configurations, specifically varying maximum deformation distances $\bar{\zeta}$. We evaluate a single H-FNO model, pre-trained on a medium deformation distance of $\bar{\zeta}=0.5\lambda$, on two distinct and challenging scenarios. The smaller range $\bar{\zeta}=0.25\lambda$ tests the model ability to interpolate, assessing its robustness when the physical diversity of the FIM is less than what it was trained on. Conversely, the larger range $\bar{\zeta}=\lambda$ tests the model ability to extrapolate, evaluating whether the learned physical principles hold for deformations beyond the scope of its training data. The results demonstrate that the model maintains excellent performance in both the interpolation and extrapolation settings. This indicates that the proposed H-FNO is not simply overfitting to a specific configuration but is learning the underlying physical relationship between deformation and the channel response.
@@ V-D Interpretability and Visualization of the Proposed H-FNO
Fig. 9 visualizes the learned weights of the first Fourier layer within the H-FNO architecture. The result shows the average magnitude of the complex-valued weights $R_{\phi}$, applied to each Fourier mode, which corresponds to a specific spatial frequency. A key observation is that the learned filter is not a simple, isotropic low-pass filter. Instead, it exhibits a more complex, anisotropic characteristic that is perfectly adapted to the 2D physical structure of the FIM array being processed as a 1D vector. The filter assigns high importance to the lowest frequency modes, e.g., modes 0-2, which are essential for capturing the large-scale, slowly varying components of the channel across the entire metasurface. However, the most prominent and insightful feature is the sharp, dominant peak at mode 8. Given that the $\times 8$ 8 FIM array is flattened column-wise, a spatial period of 8 elements corresponds to the fundamental frequency along the second spatial dimension, i.e., z-axis. The strong peak at mode 8 reveals that the model has autonomously learned that to accurately reconstruct the channel, it must prioritize not only the large-scale trends but also the distinct periodic structure occurring within each column of the FIM. This demonstrates that the proposed H-FNO learns a physically-grounded and non-trivial set of features from the data representation. Rather than applying a simple smoothing function, it has discovered the anisotropic nature of the problem and developed a specialized filter that is explicitly tailored to the unique geometry of the FIM array.
Fig. 10 compares feature activation value from a shallow high-resolution layer and the deepest low-resolution layer of H-FNO, where the feature activation refers to the numerical output of the internal network layers. For visual clarity, we provide the activation value of the first three of the feature channels, i.e., Channel 1, 2, and 3. The features from the high-resolution layer are highly oscillatory, capturing the fine-grained details and rapid phase variations of the channel between adjacent antenna elements. In contrast, the features at the low-resolution layer are significantly smoother, representing the macro-scale trends and the overall envelope of the channel response across the FIM. This result confirms that the proposed H-FNO effectively decomposes the complex channel estimation problem by learning features at multiple scales.
Figure 11: Comparison between true and predicted channel gain versus deformation distance.
@@
Fig. 11 compares the channel gain tracking performance of the proposed H-FNO and the sparsity-based OMP estimator against the true channel gain, as antenna deformation distance varies. The results first demonstrate that the true channel gain exhibits significant fluctuations as a function of the physical deformation, which is the key interference phenomenon that FIMs aim to leverage. The H-FNO prediction shows remarkable accuracy, closely tracking the true channel curve in both the less complex ($L$=5) and the more complex ($L$=20) scattering environments, validating its ability to learn a robust and continuous mapping. In contrast, while the sparsity-based OMP estimator performs reasonably well for the sparser channel with $L$=5, its performance degrades substantially in the richer scattering scenario with $L$=20. As the number of scatter paths increases, the inherent sparsity of channel is reduced, causing OMP to fail in reconstructing the correct interference pattern. This comparison highlights the superior robustness and generalization capability of the data-driven H-FNO approach, particularly in complex channel environments where rigid model assumptions may no longer hold.
Fig. 12 presents a 3D visualization of the channel gain landscape as a function of both the antenna index and the continuous deformation distance. The leftmost subfigure reveals the complex ground truth surface of the true channel, characterized by numerous peaks and troughs that demonstrate the highly non-linear variation of channel gain with respect to the FIM deformation. The predicted landscape of the proposed H-FNO remarkably reconstructs this intricate surface with high fidelity, correctly identifying the locations and magnitudes of the gain variations. Conversely, the OMP-reconstructed landscape is visibly distorted and flattened, failing to capture the deep nulls and high peaks of the true channel. This provides compelling qualitative evidence that the H-FNO has successfully learned a high-fidelity surrogate model for the true physical channel operator, accurately mapping the joint spatial-deformation domain.

### v2

@@ IV-B FIM Channel Estimation as Neural Operator Learning
A naive application of generic black-box models is ill-suited for this problem, as they lack the necessary inductive biases to handle the continuous deformation space, disregard the underlying wave physics, and suffer from mesh-dependence, making them impractical and data-inefficient [42]. In this work, we reformulate the FIM channel estimation problem as learning an operator $\mathcal{G}$ that maps the channel response function at a given set of pilot deformations $\{\bm{\zeta}_{m}\}_{m=1}^{M}$, along with a query for a target deformation, to the channel vector at that target deformation. The goal is to learn the operator $\mathcal{G}$ such that for any arbitrary and continuous target deformation $\mathbf{\bm{\zeta}}_{\text{target}}$, we have the corresponding channel estimate $\hat{\mathbf{h}}(\mathbf{\bm{\zeta}}_{\text{target}})$ as
@@ IV-B1 Lifting Layer
The input to the FNO is a function defined on the spatial domain of the FIM, which is discretized by $N$ antenna elements. For each antenna element $n$, we construct a high-dimensional feature vector $\mathbf{v}_{\text{in}}(n)\in\mathbb{R}^{3M+1}$ that encapsulates all available information: the measured channel responses and the FIM shape information, ans can be defined as
@@ IV-B2 Fourier Layers
where $\sigma$ is a non-linear activation function. The update of the Fourier layer is composed of two parallel paths: local transformation $\mathcal{W}_{t}\mathbf{v}_{t}(n)$ and global kernel integration $\left(\mathcal{K}(\mathbf{v}_{t};\mathbf{\varphi}_{t})\right)(n)$. The term $\mathcal{W}_{t}\mathbf{v}_{t}(n)$ is a pointwise linear transformation applied at each location $n$ with learnable weight $\mathcal{W}_{t}$, typically implemented as a 1D convolution with a $\times1$ kernel. It captures local and channel-wise relationships. The term $\left(\mathcal{K}(\mathbf{v}_{t};\mathbf{\varphi}_{t})\right)(n)$ is equivalent to a pointwise multiplication in the Fourier domain:
@@
Remark 3: The Fourier-domain operation allows the FNO to efficiently learn global dependencies across the entire FIM array, granting it the powerful property of mesh-independence. This distinguishes the FNO from standard convolutional networks, which learn spatially localized kernels that are inherently tied to the resolution of the training grid. The learnable weights of FNO are parameterized in the frequency domain, operating on a fixed number of low-frequency Fourier modes. Since these modes represent continuous basis functions, their meaning is independent of the underlying spatial discretization. By learning in this resolution-agnostic domain, the FNO approximates a continuous physical operator rather than a discrete pattern. Consequently, a model trained on an FIM with $N$ antennas can be directly evaluated on a larger-scale FIM with $N^{\prime}>N$ antennas. This offers immense practical value for deploying solutions across FIM hardware of varying scales and configurations. While both the proposed FNO and sparsity-based parametric channel estimation methods operate in a transformed domain related to spatial frequencies, their philosophies and implementations are fundamentally different. The sparsity-based methods solve a sparse linear inverse problem on a discrete grid, while FNO learns a continuous and non-linear operator that is biased towards smooth functions.
@@ IV-B3 Projection Layer
$\begin{split}\left[\Re\left\{\hat{\mathbf{h}}(\mathbf{\bm{\zeta}}_{\text{target}})_{n}\right\},\Im\left\{\hat{\mathbf{h}}(\mathbf{\bm{\zeta}}_{\text{target}})_{n}\right\}\right]^{T}=\mathcal{Q}\left(\mathbf{v}_{T}(n)\right).\end{split}$ (48)
@@ IV-C1 Encoder Path
The output is then downsampled by an average pooling operator $\mathcal{A}$, i.e., $\mathbf{v}_{l}=\mathcal{A}(\mathbf{v}^{\prime}_{l})$. This operation halves the spatial resolution of the feature map while keeping its channel dimension $d_{v}$ constant. This process is repeated for $L_{\text{enc}}$ levels, resulting in a sequence of latent representations $\{\mathbf{v}^{\prime}_{1},\mathbf{v}^{\prime}_{2},\dots,\mathbf{v}^{\prime}_{L_{\text{enc}}}\}$ with progressively decreasing spatial dimensions ($N,N/2,\dots,N/2^{L_{\text{enc}}-1}$).
@@ IV-C3 Decoder Path
where $\left[\cdot;\cdot\right]$ denotes concatenation along the channel dimension. Specifically, the upsampled feature map with $d_{v}$ channels is combined with the feature map from the encoder path, resulting in a fused feature map with a doubled channel dimension of $2d_{v}$. Finally, another FNO block is applied to fuse these multi-scale features $\mathbf{u}_{l}=\mathcal{B}_{\text{FNO}}(\mathbf{u}_{l,\text{cat}})$. After $L_{\text{enc}}$ decoding steps, the final high-resolution latent representation is passed through the projection layer $\mathcal{Q}$ to produce the channel estimate.
@@
As illustrated in Section VI-A, the proposed H-FNO $\mathcal{G}_{\Theta}$ can be interpreted as a learned approximation to the MAP estimator for the FIM channel estimation problem, where the underlying prior imposed by the H-FNO architecture aligns with the inherent spectral properties of the physical channel model. The mechanism of truncating high-frequency Fourier modes in Eqs. (45)-(47) imposes a strong spectral bias on the solution. This architectural choice acts as a powerful structural regularizer, forcing the learned channel function to be inherently smooth and band-limited. This approach is conceptually analogous to the role of the prior in a MAP framework, as both methods effectively constrain the solution to a physically plausible function space. Specifically, the GP is a collection of random variables, any finite number of which have a joint Gaussian distribution [44]. In this paper, we regard the channel function ${h}(\bm{\zeta})$ as a GP, mapping deformation $\bm{\zeta}$ to channel response. The FNO truncation of high-frequency Fourier modes implements a smoothness constraint, effectively acting as a hard spectral prior, whereas the GP uses a soft probabilistic prior via the kernel. Both approaches favor band-limited channel functions, but the FNO achieves this through architectural design rather than probabilistic modeling.
@@
For the data fidelity term of MAP, i.e., the first term in Eq. (40), the H-FNO is trained end-to-end by minimizing the MSE loss in Eq. (52), which is a direct data-driven approximation of the data fidelity term. For the second regularization term in Eq. (40) that penalizes high-frequency channel content, the H-FNO implements this regularization structurally. Its architectural truncation of Fourier modes beyond $\kappa_{\text{max}}$ is not merely a soft penalty but a hard constraint. This is structurally equivalent to imposing a MAP prior $p(\bar{\mathbf{h}})$, where the probability of any function containing these higher frequencies is explicitly zero. Consequently, the network is architecturally incapable of representing non-band-limited functions, forcing the training process to find a solution entirely within a physically-motivated function space that adheres to this strong spectral bias. Within the preserved modes, the learnable transformation $\mathbf{R}_{\varphi}$ can be interpreted as learning the inherent structure of the channel function space.
@@
Remark 4: In the popular deep learning-enabled channel estimators, the deep neural network does not require an explicit prior, which implicitly imposes a learned structural prior on the function space of possible channels as a generic function approximator. However, the proposed H-FNO is not a closed box but a principled and data-driven method for finding a powerful approximation $\mathcal{G}_{\Theta}$, to the true MAP estimator. The FNO architecture is explicitly designed to search within a function space constrained by a physically-motivated spectral bias, which is structurally enforced by truncating high-frequency components in the Fourier domain. This theoretical synergy between the inductive biases of the proposed H-FNO and the physical characteristics of FIM provides a rigorous justification for its application.
IV-D Complexity Analysis
IV-D1 Computational Complexity
In the proposed H-FNO architecture, the primary computational bottleneck is the FFT operation within each Fourier layer. For an FIM with $N$ radiating elements and a latent channel dimension of $d_{v}$, a single FFT has a complexity of $\mathcal{O}(N\log N)$. While the H-FNO processes features across multiple spatial resolutions, the overall cost is dominated by the layers with the highest resolution. Therefore, the total inference complexity is approximately $\mathcal{O}(Ld_{v}N\log N)$, where $L=2{L_{\text{enc}}}+1$ is the total number of Fourier layers in the network. This demonstrates that the complexity scales quasi-linearly with the FIM size, making the H-FNO efficient even for large arrays. The training complexity of the proposed H-FNO architecture for a single epoch depends on the cost of the forward and backward passes, repeated for all batches. As the complexity of backpropagation is a small constant multiple of the forward pass, the total training complexity per epoch is given by $\mathcal{O}\left(\frac{N_{s}}{B_{s}}Ld_{v}N\log N\right)$, where $N_{s}$ is the total number of training samples and $B_{s}$ is the batch size in the training stage.
IV-D2 Sample Complexity
For the considered channel estimation problem in FIM systems, the sample complexity, i.e., the number of pilot measurements $M$ required for generalization, is determined by the intrinsic complexity of the channel function $\mathbf{h}(\bm{\zeta})$. This complexity relates to the function smoothness in the deformation space. Following principles analogous to the Nyquist-Shannon theorem, a smoother function requires fewer measurements to be accurately characterized. This concept is directly embodied in the H-FNO design. The number of retained Fourier modes $\kappa_{\text{max}}$, acts as an architectural prior on the smoothness of the learned operator. A smaller $\kappa_{\text{max}}$ constrains the model to represent smoother functions, thereby aligning the model’s capacity with the information available from a limited number of pilot measurements $M$.
@@ V Numerical Results
In this section, we provide extensive numerical results to validate the effectiveness of the proposed H-FNO framework by comparing it against several model-based benchmarks, and also present the interpretability analysis of H-FNO.
V-A Parameter Setups and Dataset Construction
[…]
In the training stage of the proposed H-FNO, a dataset of $N_{s}=20,000$ samples is synthetically generated, with $N_{v}=1000$ of these samples being reserved for the validation set. For each sample, a unique channel environment is realized based on the clustered multipath model, and then $M$ noisy pilot observations are generated, where the corresponding noise-free channel vector $\mathbf{h}_{k}(\bm{\zeta}_{\text{target}})$ serves as the ground-truth label. To enhance the model robustness, the SNR for each training batch is randomly selected from {0, …, 20} dB. The model is trained for $E_{s}=50$ epochs with a batch size of $B_{s}=64$. For each SNR point under performance testing, the NMSE results are averaged over $U=500$ Monte Carlo simulations. In each simulation trial, both a new channel realization and a new random target deformation shape are generated to ensure a robust and fair evaluation. For the benchmark algorithms, the number of nearest neighbors for KNN is set to $\bar{k}=5$, and the OMP dictionary is constructed from a discrete grid of $D=256$ angles. For the SBL estimator, the iterative learning process is controlled by a maximum of $100$ iterations and a convergence tolerance of $10^{-4}$. For the H-FNO, we set the latent dimension to $d_{v}=64$ and the number of encoder levels to $L_{\text{enc}}=2$, with the number of Fourier modes per level set as $\kappa_{l,\text{max}}=16/2^{l}$, which represent a standard architecture designed to balance model expressiveness with computational efficiency.
@@ V-B Performance Comparison between Different Algorithms
Fig. 5 presents the NMSE performance of different channel estimation algorithms versus the signal-to-noise ratio (SNR), where the channel estimation accuracy of all methods improves with increasing SNR. Among the conventional model-based methods, the parametric channel estimation schemes, i.e., OMP and SBL, show superior channel estimation accuracy, benefiting from the inherent sparsity of the mmWave channel, while interpolation-based methods provide a stable but less accurate estimation. In particular, the SBL algorithm can offer superior performance compared to OMP by avoiding local minima and providing a probabilistic estimate of the channel parameters. The proposed H-FNO model consistently and significantly outperforms all other conventional model-based algorithms across the entire SNR range, demonstrating its powerful capability in learning the complex mapping from FIM shapes to channel responses and its robustness against noise. Note that while the numerical results show the superior accuracy of a learning-based approach, the conventional model-based frameworks retain significant practical value. They are highly interpretable, provide theoretical performance guarantees, and can be deployed without large training datasets or complex offline training pipelines.
@@
Figure 7: Sum-rate performance against SNR for different channel estimation algorithms.
Fig. 7 illustrates the sum-rate performance achieved when using the estimated channels from different algorithms for zero-forcing beamforming [10]. We can observe that the proposed H-FNO framework consistently outperforms all model-based benchmarks, achieving a sum-rate significantly closer to the theoretical upper bound provided by perfect CSI. This demonstrates that the superior channel estimation accuracy of H-FNO, as demonstrated by its lower NMSE, directly translates into substantial gains in the overall spectral efficiency of FIM assisted multi-user systems.
@@ V-C Generalization Performance of the Proposed H-FNO
Figure 8: NMSE performance of different number of antennas NN for the proposed H-FNO.
@@
Fig. 9 validates the zero-shot generalization capability of the proposed H-FNO model across different FIM hardware configurations, specifically varying maximum deformation distances $\bar{\zeta}$. We evaluate a single H-FNO model, pre-trained on a medium deformation distance of $\bar{\zeta}=0.50\lambda$, on two distinct and challenging scenarios. The smaller range $\bar{\zeta}=0.25\lambda$ tests the model ability to interpolate, assessing its robustness when the physical diversity of the FIM is less than what it was trained on. Conversely, the larger range $\bar{\zeta}=1.00\lambda$ tests the model ability to extrapolate, evaluating whether the learned physical principles hold for deformations beyond the scope of its training data. The results demonstrate that the model maintains excellent performance in both the interpolation and extrapolation settings. This indicates that the proposed H-FNO is not simply overfitting to a specific configuration but is learning the underlying physical relationship between deformation and the channel response.
Figure 10: Robustness analysis under impairments and distribution shift for the proposed H-FNO.
Fig. 10 evaluates the robustness of the proposed H-FNO framework, where three types of hardware imperfections are introduced during the testing phase: 1) a pilot shape actuation error, modeled by a Gaussian perturbation with a standard deviation set to 5% of the maximum deformation range $\sigma_{\text{act}}=0.05\bar{\zeta}$; 2) oscillator phase noise, introduced as a random phase shift with a standard deviation of 0.05 radians; and 3) a channel distribution shift, where the model, trained on channels with $L=5$ clusters, is tested on a more complex rich-scattering environment with $L=20$ clusters. The results show only a slight performance degradation in the presence of the hardware impairments, demonstrating strong robustness. Furthermore, the model maintains highly effective estimation even under the significant channel distribution shift, which confirms its ability to learn generalizable physical principles rather than merely overfitting to the training statistics.
@@ V-D Interpretability and Visualization of the Proposed H-FNO
Fig. 11 visualizes the learned weights of the first Fourier layer within the H-FNO architecture. The result shows the average magnitude of the complex-valued weights $R_{\phi}$, applied to each Fourier mode, which corresponds to a specific spatial frequency. A key observation is that the learned filter is not a simple, isotropic low-pass filter. Instead, it exhibits a more complex, anisotropic characteristic that is perfectly adapted to the 2D physical structure of the FIM array being processed as a 1D vector. The filter assigns high importance to the lowest frequency modes (e.g., modes 0-2), which are essential for capturing the large-scale, slowly varying components of the channel across the entire metasurface. However, the most prominent and insightful feature is the sharp, dominant peak at mode 8. Given that the $8\times 8$ FIM array is flattened column-wise, a spatial period of 8 elements corresponds to the fundamental frequency along the second spatial dimension, i.e., the z-axis. The strong peak at mode 8 reveals that the model has autonomously learned that to accurately reconstruct the channel, it must prioritize not only the large-scale trends but also the distinct periodic structure occurring within each column of the FIM. This demonstrates that the proposed H-FNO learns a physically-grounded and non-trivial set of features from the data representation. Rather than applying a simple smoothing function, it has discovered the anisotropic nature of the problem and developed a specialized filter that is explicitly tailored to the unique geometry of the FIM array.
Fig. 12 provides insight into the multi-scale feature learning enabled by the hierarchical architecture of H-FNO. It compares feature activations from a shallow high-resolution layer (Encoder 1 of H-FNO, NN=64) with those from the deepest low-resolution layer (Bottleneck of H-FNO, $N/4=16$). The features from the high-resolution encoder are highly oscillatory, capturing the fine-grained details and rapid phase variations of the channel between adjacent antenna elements. In contrast, the features at the bottleneck are significantly smoother, representing the macro-scale trends and the overall envelope of the channel response across the FIM. This result confirms that the H-FNO effectively decomposes the complex estimation problem by learning features at multiple scales
Figure 13: Spatial error distribution of per-antenna channel estimation NMSE.
@@
Fig. 13 presents a spatial diagnostic of the estimation performance by visualizing the per-antenna NMSE across the $8\times 8$ FIM at an SNR of 10 dB. It is observed that while both estimators exhibit higher errors towards the edges of the array, i.e., a common artifact in array processing, the H-FNO demonstrates significantly superior performance. Compared to the SBL algorithm, the proposed H-FNO features a substantially larger dark region, indicating a lower error floor across the central elements. This illustrates that the H-FNO not only achieves a lower overall NMSE but also provides a more uniformly accurate channel estimate across the entire FIM, which is crucial for subsequent signal processing operations.
Fig. 14 presents a 3D visualization of the channel gain landscape as a function of both the antenna index and the continuous deformation distance, where the channel gain represents the magnitude of the channel response at each individual radiating element, expressed in decibels as $20\log_{10}(|\mathbf{h}_{{{\zeta}_{n}},n}|)$ (dB). The leftmost subfigure reveals the complex ground truth surface of the true channel, characterized by numerous peaks and troughs that represent the optimal deformation positions for each antenna. The predicted landscape of H-FNO remarkably reconstructs this intricate surface with high fidelity, correctly identifying the locations and magnitudes of the gain variations. Conversely, the OMP-reconstructed landscape is visibly distorted and flattened, failing to capture the deep nulls and high peaks of the true channel. This provides compelling qualitative evidence that the H-FNO has successfully learned a high-fidelity surrogate model for the true physical channel operator, accurately mapping the joint spatial-deformation domain.